R1 = 4.5 ohms, R2 = 3.6 ohms, and R3 = 2.8 ohms

RI R2

A) What is the equivalent resistance?

B) If these resistors are connected across a 12V battery, what's the current through R3?
Answer #1

A)

R1 and R2 are in parallel combination, hence their equivalent resistance is R12,

1/R12=1/R1+1/R2

1/R12=1/4.5+1/3.6

R12= 2.0 ohm

Now R3 is in series combination with R12. Total equivalent resistance of all three=R123

Req=R123 = R12 + R3 = 2.0 + 2.8 = 4.8 ohm

B)

I = V/Req = 12/4.8 = 2.5 A